# FRoundation: Are Foundation Models Ready for Face Recognition?

### v2

@@ 4 Experimental setups
To explore the capabilities of foundation models for face recognition, we assess their performance using the following benchmarks: Labeled Faces in the Wild (LFW) [25], Celebrities in Frontal-Profile in the Wild (CFP-FP) [37], AgeDB30 [32], Cross-age LFW (CA-LFW) [52], CrossPose LFW (CP-LFW) [51]. The results are reported on each benchmark as verification accuracies in (%) and following their official evaluation protocol. In addition, we evaluated on large-scale evaluation benchmarks, IARPA Janus Benchmark-B (IJB-B) [46], and IARPA Janus Benchmark–C (IJB-C) [31]. For IJB-C and IJB-B, we used the official 1:1 mixed verification protocol and reported the verification performance as true acceptance rates (TAR) at false acceptance rates (FAR) of 1e-3, 1e-4, and 1e-5. All these benchmarks were chosen as they are widely used to benchmark the latest FR developments and provide a comprehensive variation of use-cases [17, 42, 4, 16].
@@
To fine-tune the considered foundation models, we employ the CASIA WebFace [47] dataset, consisting of 0.5 million images and 10k identities. We also conduct large-scale fine-tuning on the MS1MV2 [17] and WebFace4M [55] datasets. MS1MV2 is a refined version of MS-Celeb-1M [22] by [17], containing 5.8M images of 85K identities. WebFace4M is a subset of the WebFace260M dataset [55], consisting of 200k identities and 4 million images. The images in the training and testing dataset are aligned and cropped to as described in [17] using five landmark points extracted by the Multi-task Cascaded Convolutional Networks (MTCNN) [50].
@@
DINOv2 [33] officially released four ViT architectures, small (Vit-S), base (Vit-B), large (Vit-L), and giant (Vit-G). ViT-S, Vit-B, Vit-L and Vit-G contain 21M, 86M, 0.3B and 1.1B parameters, respectively. All models use a patch size of 14, but they differ in embedding dimensions and the number of attention heads. On the other hand, CLIP [35] offers 4 different models with 2 architectures: base and large. The base model Vit-B of CLIP contains 86M parameters and has 2 variants with different patch sizes, 16 and 32. The large model ViT-L/14 has 0.3 billion parameters and includes a variant that was pre-trained at a higher resolution of 336 pixels for one additional epoch to boost performance [40]. We will start by evaluating the performance of all models in Section 5.1. For further detailed experiments, we focus on the following models: ViT-S for DINOv2 and ViT-B for CLIP, as these are the smallest models released by the corresponding authors, which makes our detailed experiments viable (given hardware and time limitations). We also investigate the performance of the larger models in Section 5.2 and choose ViT-L for CLIP and DINOv2 for a fairer comparison, as the largest model, namely the giant model of DINOv2, has 1.1 billion parameters compared to 0.3 billion parameters for DINOv2 and CLIP ViT Large.
@@
We utilize the CosFace [42] loss function to train all the presented models in this paper with a margin penalty of and scale factor of , following [42], as well as other works analyzing different building blocks of face recognition [5]. During the fine-tuning, we used AdamW [30] optimizer with a weight decay of 0.05. We train for 40 epochs (on CaisaWebFace) and for 30 epochs (on MS1MV2 and WebFace4M), with a batch size set to 512 [17]. The initial learning rate is set to 0.0001 and is updated using a Cosine Learning Rate scheduler [29]. Additionally, the LoRA rank is set to 16 for all applicable experiments. The images are resized to pixels to adapt to the image resolution initially used by DINOv2 and CLIP during training. For data augmentation, we apply horizontal flipping and RandAug [14] with a number of operations and a magnitude of , following [6]. We additionally trained 12 instances of Vit, Vit-S and Vit-L, from scratch on different subsets of CASIA-WebFace as well as on MS1MV2 and WebFace4M. These models are noted as Baseline.
| # Identities | # Images | Method | LFW | CFPFP | AgeDB30 | CALFW | CPLFW | Avg. | IJB-B | IJB-C | ||||
@@ 5.1 Are Foundation Models ready for Face Recognition?
We first evaluate the face verification performances of the considered foundation models, DINOv2 and CLIP, on the several challenging benchmarks described in Section 4. For DINOv2 and CLIP, we utilized the official pretrained models released by the corresponding authors [33, 35]. The models, in this case, are utilized as feature extractors without fine-tuning the model weights.
The achieved results by the different pretrained architectures of DINOv2 and CLIP are presented in Table 1. All network architectures are based on vision-transformer (Vit). For details on network architectures, one can refer to the corresponding papers [33, 35]. The results of the first row (noted as Baseline (MS1MV2)) are achieved by a Vit-S model trained from scratch on MS1MV2. These evaluation results are provided in this table as a reference. One can clearly observe that different pretrained CLIP model versions outperformed DINOv2 models on the considered benchmarks. This result might be attributed to the fact that the curated dataset LVD-142M [33], on which DINOv2 was trained, involved a post-processing step that blurred identifiable faces [33]. Although the considered foundation models are not trained and optimized to perform feature extraction for face verification, they achieved relatively high accuracies on considered benchmarks. For example, the achieved verification accuracies by CLIP (ViT-L/14@336), the largest model from CLIP, was on LFW. On benchmarks with cross-age evaluation protocol, the best-achieved accuracies on AgeDB30 and CALLFW were and , respectively. All considered models achieved relatively low TAR at the reported thresholds of FAR on large and challenging benchmarks, IJB-B and IJB-C.
From the reported results in Table 1, one can conclude that the achieved results by foundation models on face verification are far from being random. They achieved relatively high accuracies on less challenging benchmarks (LFW). However, this performance significantly drops when the evaluation benchmark contains challenging pairs such as AgeDB30, CALFW, IJB-B and IJB-C, especially when considering the baseline model trained from scratch and the current performances of state-of-the-art FR solutions [17, 42, 4, 16].
@@ 5.2 FRoundation: Fine-tuning Foundation Models
Table 2 presents the verification accuracies achieved by considered foundation models, DINOv2 and CLIP, fine-tuned with LoRA (Section 3.3) on different subsets of CASIA-WebFace. The results noted with baseline are achieved by ViT-S trained from scratch and provided as a reference. In this experiment, we utilized the smallest released pretrained ViT architecture from DINOv2 (ViT-S) and CLIP (ViT-B). The results of the first two rows in Table 2 are achieved by pretrained DINOv2 (ViT-S) and CLIP (ViT-B) without finetuning. Driven by possible technical limitations and ethical concerns in practice to collect large-scale face (or other biometric) datasets [7], we propose to fine-tune the foundation models on small subsets from CAISA-WebFace and compare them to the case where the model was trained from scratch. Specifically, we provided a border evaluation of the impact of dataset size in terms of number of identities (dataset width) on the model performance by utilizing (n)K identities from CAISA-WebFace, where . Additionally, we investigate the impact of the dataset depth (number of images) on the model performance by comparing the case where the 2.5k identities are randomly selected to the case where these 2.5K identities are selected (noted as 2.5k (S)) with the largest number of images per identity. The 2.5K and 2.5k (S) resulted in a total of 163945 and 289228 images, respectively. Based on the reported results in Table 2, we made the following observations:
@@
Fine-tuning DINOv2 and CLIP on different subsets from CASIA-WebFace significantly improved the achieved verification accuracies on all evaluation benchmarks, in comparison to the case where DINOv2 and CLIP are utilized without fine-tuning. Initially, without fine-tuning, DINOv2 (ViT-S) and CLIP (ViT-B) achieved average accuracies of 64.70 and 82.64, respectively. These average accuracies increased to 87.10 and 90.75 by fine-tuning DINOv2 and CLIP, respectively on a small subset of only 1k identities from CAISA-WebFace.
@@
As expected, using a larger subset for fine-tuning DINOv2 and CLIP consistently resulted in higher verification accuracies across all considered benchmarks. The same observation can be also made for the model trained from scratch (Baseline).
@@
Finetuning ViT of pretrained DINOv2 and CLIP led to higher verification accuracies, in comparison to the case where the model is trained from scratch. The superiority of foundation models over the model trained from scratch on CASIA-WebFace (or a subset of CASIA-WebFace) can be observed on the small evaluation benchmarks and on the large-scale, IJB-B and IJB-C, benchmarks. This observation can be seen also in Figure 3.
@@
Figure 3 presents average accuracies (y-axis) of baseline and foundation models fine-tuned on different subsets (x-axis) from CASIA-WebFace. One can notice that the average accuracies slightly improved when a larger subset of CASIA-WebFace is utilized. On the other hand, the performance of the baseline significantly increases with more data, starting at an average accuracy of 69.96% when trained on 1K and reaching 88.56% on the full dataset with an increase of 18.6%. The same observation can be also made on IJB-B and IJB-C benchmarks (Table 2).
@@
The impact of the dataset depth (number of images) can be observed by comparing the achieved results of models trained/finetuned on 2.5K (163945) and 2.5K (S) (289228). One can notice that increasing the dataset depth leads to higher verification accuracies in most of the settings.
@@ 5.3 Large-scale Fine-tuning
Table 3 presented the achieved recognition performances by models trained from scratch (baseline) and fine-tuned (CLIP and DINOV2) on a relatively small dataset, CASIA-WebFace (0.5m images) and larger datasets, MS1MV2 (5.8m images) and WEbFace4M (4m images). All results are reported using large (Vit-L) and small (Vit-S of DINOv2 and baseline and Vit-B of CLIP) architectures. We made the following observation from the reported results in Table 3:
Using CASIA-WebFace (0.5m images), finetuning DINOv2 (Vit-S and Vit-L architectures) and CLIP (Vit-B and Vit-L architectures) led to higher verification accuracies, in comparison to the case where Vit-S and Vit-L are trained from scratch on the same data. For example, the average verification accuracies on the small benchmarks, LFW, CFPFP, AgeDB30, CALFW, and CPLFW, was 87.39% by baseline (Vit-L) which is lower than 94.28% and 94.26% achieved by the fine tuned DINOv2 and CLIP, respectively.
The models that are trained/finetuned on large datasets, MS1MV2 (5.8m images) and WEbFace4M (4m images), achieved higher verification accuracies than the one trained on the relatively smaller dataset, CASIA-WebFace (0.5m images).
Using small architectures (Vit-S and Vit-B) and large-scale training datasets (MS1MV2 or WebFace4M), the model trained from scratch achieved higher verification accuracies than the finetuned DINOv2 and CLIP. This observation can be noticed from the achieved results on small benchmarks as well as on large-scale, IJB-B and IJB-C, benchmarks.
Using large architectures (Vit-L) and large-scale training/finetuning datasets (MS1MV2 or WebFace4M), the finetuned DINOv2 and CLIP achieved slightly higher recognition performance than the models trained from scratch on most of the evaluation benchmarks.
To conclude, for the case where only a relatively small training dataset is accessible, finetuning pretrained foundation models can achieve higher recognition accuracy than training the same model architecture from scratch. In the case that one has access to large training datasets, training a model from scratch for face recognition can achieve very competitive results to finetuning foundation models. However, this, next to the technical and legal limitations of collecting or maintaining the data, comes with a high training time cost. For example, training Vit-L from scratch on MS1MV2 using the settings described in Section 4 requires around 70 GPU hours on MS1MV2, in comparison to around 40 GPU hours for finetuning DINOv2 with LoRA on 8 Nvidia A100 SXM4 40 GB GPUs.
| Method | Backbone | Train data | LFW | CFPFP | AgeDB30 | CALFW | CPLFW | Avg. | IJB-B | IJB-C | ||||

### v3

@@ 4 Experimental setups
To explore the capabilities of foundation models for FR, we assess their performance using the following benchmarks: Labeled Faces in the Wild (LFW) [31], Celebrities in Frontal-Profile in the Wild (CFP-FP) [54], AgeDB30 [46], Cross-age LFW (CA-LFW) [72], CrossPose LFW (CP-LFW) [71]. The results are reported on each benchmark as verification accuracies in (%) and following their official evaluation protocol. In addition, we evaluated on large-scale evaluation benchmarks, IARPA Janus Benchmark-B (IJB-B) [64], and IARPA Janus Benchmark–C (IJB-C) [43]. For IJB-C and IJB-B, we used the official 1:1 mixed verification protocol and reported the verification performance as true acceptance rates (TAR) at false acceptance rates (FAR) of 1e-3, 1e-4, and 1e-5. All these benchmarks were chosen as they are widely used to benchmark the latest FR developments and provide a comprehensive variation of use-cases [20, 60, 5, 19]. All results in the paper are reported using cross-validation settings, ensuring no overlap between the training and testing datasets. This approach aligns with SOTA works [20, 5, 45] in FR, by utilizing datasets such as CASIA-WebFace, MS1MV2, and WebFace4M for training/finetuning while reporting verification results on mainstream evaluation benchmarks, including LFW, CFP-FP, AgeDB30, CALFW, CPLFW, CFP-FP, IJB-C and IJB-B.
@@
To fine-tune the considered foundation models, we employ the CASIA-WebFace [66] dataset, consisting of 0.5 million images and 10K identities. We also conduct large-scale fine-tuning on the MS1MV2 [20] and WebFace4M [75] datasets. MS1MV2 is a refined version of MS-Celeb-1M [27] by [20], containing 5.8M images of 85K identities. WebFace4M is a subset of the WebFace260M dataset [75], consisting of 200K identities and 4 million images. We also investigate the performance of the selected foundation models when trained using synthetic data, focusing on a latent diffusion models-based approach, IDiff-Face [6], and a generative adversarial network (GAN)-based approach, SFace2 [7]. For both synthetic datasets, we conducted training on 10K identities, using 50 images per identity. The images in the training and testing dataset are aligned and cropped to as described in [20] using five landmark points extracted by the Multi-task Cascaded Convolutional Networks (MTCNN) [70].
@@
DINOv2 [47] officially released four ViT architectures, small (ViT-S), base (ViT-B), large (ViT-L), and giant (ViT-G). ViT-S, ViT-B, ViT-L, and ViT-G contain 22M, 86M, 0.3B, and 1.1B parameters, respectively. All models use a patch size of 14 but differ in embedding dimensions and the number of attention heads, with ViT-S having an embedding dimension of 384 and 6 heads, ViT-B having an embedding dimension of 768 and 12 heads, ViT-L having an embedding dimension of 1024 and 16 heads, and ViT-g having an embedding dimension of 1536 and 24 heads. On the other hand, CLIP [51] offers 4 different models with 2 architectures: base and large. The base model ViT-B of CLIP contains 86M parameters and has 2 variants with different patch sizes, 16 (ViT-B/16) and 32 (ViT-B/32). The large model ViT-L/14 has 0.3 billion parameters and includes a variant, namely ViT-L/14@336, that was pre-trained at a higher resolution of 336 pixels for one additional epoch to boost performance [58]. The ViT-B model has a width of 768 and 12 attention heads, while the ViT-L model has a width of 1024 and 16 attention heads. We will start by evaluating the performance of all models in Section 5.1. For further detailed experiments, we focus on the following models: ViT-S for DINOv2 and ViT-B/16 for CLIP, as these are the smallest models released by the corresponding authors, which makes our detailed experiments viable (given hardware and time limitations). We also investigate the performance of the larger models in Section 5.2 and choose ViT-L for CLIP and DINOv2 for a fairer comparison, as the largest model, namely the giant model of DINOv2, has 1.1 billion parameters compared to 0.3 billion parameters for DINOv2 and CLIP ViT Large.
@@
We utilize the CosFace [60] loss function to train all models presented in this paper with a margin penalty of and scale factor of , following [60], as well as other works analyzing different building blocks of FR [7]. During the fine-tuning, we used AdamW [42] optimizer with a weight decay of . We train for epochs (on CASIA-WebFace) and for epochs (on MS1MV2 and WebFace4M), with a batch size set to [20]. The initial learning rate is set to and is updated using a Cosine Learning Rate scheduler [41]. Additionally, the LoRA rank is set to for all applicable experiments. The images are resized to pixels to adapt to the image resolution initially used by DINOv2 and CLIP during training. For data augmentation, we apply horizontal flipping and RandAug [17] with operations and a magnitude of , following [8]. During training and following [20, 5], we explore efficient face verification datasets (e.g. LFW, CALFW, CPLFW, CFP-FP, AgeDB) to track and check the convergence status of the model. This has been performed after each epoch. It is important to note that for the RFW [62] evaluation, models trained on MS1MV2 are not subject to full cross-validation due to the identity overlap between RFW and MS1MV2, which is a cleaned version derived from MS-Celeb-1M [27], as stated in [62]. We additionally trained instances of ViT, ViT-S, and ViT-L, from scratch on different subsets of CASIA-WebFace as well as on MS1MV2 and WebFace4M. These models are noted as the baseline.
Computational Cost
We utilized two foundation models in this paper, each with different base architectures. For DINOv2, we utilized ViT-S and ViT-L, containing 22M and 0.3B parameters, respectively. For CLIP, we utilized ViT-B and ViT-L, containing 86M and 0.3B parameters, respectively. Given that these models utilize 32 floating-point (4-byte) values to represent each parameter, the required memory footprint for each model is 4 times the number of parameters. To evaluate computational time during inference, we measure the models’ latency on a Quadro RTX 6000 GPU. Specifically, the ViT-S and ViT-L variants of DINOv2 require 6.993 ms and 13.614 ms, respectively, to process a single image on this hardware. When fine-tuning the models, we attach LoRA layers to the various considered model architectures. These layers introduce computational overhead, as the forward pass of the model becomes more complex due to the need to process the additional weights. As a result, the forward pass time for the DINOv2 ViT-S and ViT-L increases to 9.862 ms and 19.502 ms, respectively. For the CLIP variants, namely ViT-B and ViT-L, they require 8.657 ms and 12.74 ms, respectively, which increase to 12.498 and 18.817 ms when LoRA layers are added. While models fine-tuned with LoRA exhibit a slower forward pass during training, they benefit from the fact that only the additional LoRA layers are updated during the backpropagation process. For DINOv2, ViT-S (22M parameters) and ViT-L (0.3B parameters) incorporate an additional 0.3M and 1.5M parameters due to LoRA, respectively, which corresponds to only 1.32% and 0.51% of the total trainable parameters. For CLIP, ViT-B (86M parameters) and ViT-L (0.3B parameters) also add 0.5M and 1.5M parameters from LoRA, respectively, which account for only 0.68% and 0.51% of the total trainable parameters. It is important to note that LoRA adapter weights are merged with the base model weights once training is complete, as discussed in Section 3.2. This merging reduces memory usage and allows for inference speeds that are comparable to those of the original model, eliminating the need to maintain separate layers [30]. As a result, the extra computational load and memory usage during training is removed during inference.
| # Identities | # Images | Method | LFW | CFP-FP | AgeDB30 | CALFW | CPLFW | Avg. | IJB-B | IJB-C | ||||
@@ 5.1 Are Foundation Models ready for Face Recognition?
We first evaluate the face verification performances of the considered foundation models, DINOv2 and CLIP, on several challenging benchmarks described in Section 4. For DINOv2 and CLIP, we utilized the official pre-trained models released by the corresponding authors [47, 51]. The models, in this case, are utilized as feature extractors without fine-tuning the model weights.
The results achieved by the different pre-trained architectures of DINOv2 and CLIP are presented in Table 1. All network architectures are based on vision-transformer (ViT). For details on network architectures, one can refer to the corresponding papers [47, 51]. The results of the first row (noted as baseline (MS1MV2)) are achieved by a ViT-S model trained from scratch on MS1MV2. These evaluation results are provided in this table as a reference. One can clearly observe that different pre-trained CLIP model versions outperformed DINOv2 models on the considered benchmarks. This result might be attributed to the fact that the curated dataset LVD-142M [47], on which DINOv2 was trained, involved a post-processing step that blurred identifiable faces [47]. Although the considered foundation models are not trained and optimized to perform feature extraction for face verification, they achieved relatively high accuracies on the considered benchmarks. For example, the achieved verification accuracy by CLIP (ViT-L/14@336), the largest model from CLIP, was on LFW. On the benchmarks with cross-age evaluation protocol, the best-achieved accuracies on AgeDB30 and CALLFW were and , respectively. All considered models achieved relatively low TAR at the reported thresholds of FAR on large and challenging benchmarks, IJB-B and IJB-C.
From the reported results in Table 1, one can conclude that the achieved results by foundation models on face verification are far from being random. They achieved relatively high accuracies on less challenging benchmarks (LFW). However, this performance significantly drops when the evaluation benchmark contains challenging pairs such as AgeDB30, CALFW, IJB-B, and IJB-C, especially when considering the baseline model trained from scratch and the current performances of SOTA FR solutions [20, 60, 5, 19].
@@ 5.2 FRoundation: Fine-tuning Foundation Models
Table 2 presents the verification accuracies achieved by considered foundation models, DINOv2 and CLIP, fine-tuned with LoRA (Section 3.3) on different subsets of CASIA-WebFace. The results noted as the baseline are achieved by ViT-S trained from scratch and provided as a reference. In this experiment, we utilized the smallest released pre-trained ViT architecture from DINOv2 (ViT-S) and CLIP (ViT-B). The results of the first two rows in Table 2 are achieved by pre-trained DINOv2 (ViT-S) and CLIP (ViT-B) without fine-tuning. Driven by possible technical limitations and ethical concerns in practice to collect large-scale face (or other biometric) datasets [9], we propose to fine-tune the foundation models on small subsets from CAISA-WebFace and compare them to the case where the model was trained from scratch. Specifically, we provided a border evaluation of the impact of dataset size in terms of number of identities (dataset width) on the model performance by utilizing K identities from CASIA-WebFace, where . Additionally, we investigate the impact of the dataset depth (number of images) on the model performance by comparing the case where the K identities are randomly selected to the case where these identities are selected (noted as K(S)) with the largest number of images per identity. The K and K(S) resulted in a total of and images, respectively. Based on the reported results in Table 2, we made the following observations:
@@
Fine-tuning DINOv2 and CLIP on different subsets from CASIA-WebFace significantly improved the achieved verification accuracies on all evaluation benchmarks, in comparison to the case where DINOv2 and CLIP are utilized without fine-tuning. Initially, without fine-tuning, DINOv2 (ViT-S) and CLIP (ViT-B) achieved average accuracies of % and %, respectively. These average accuracies increased to and by fine-tuning DINOv2 and CLIP, respectively, on a small subset of only 1K identities from CAISA-WebFace.
@@
As expected, using a larger subset for fine-tuning DINOv2 and CLIP consistently resulted in higher verification accuracies across all considered benchmarks. The same observation can also be made for the model trained from scratch (baseline).
@@
Fine-tuning ViT of pre-trained DINOv2 and CLIP led to higher verification accuracies, in comparison to the case where the model is trained from scratch. The superiority of foundation models over the model trained from scratch on CASIA-WebFace (or a subset of CASIA-WebFace) can be observed on the small evaluation benchmarks and the large-scale, IJB-B and IJB-C, benchmarks. This observation can be seen also in Figure 3.
@@
Figure 3 presents average accuracies (y-axis) of baseline and foundation models fine-tuned on different subsets (x-axis) from CASIA-WebFace. One can notice that the average accuracies slightly improved when a larger subset of CASIA-WebFace is utilized. On the other hand, the performance of the baseline significantly increases with more data, starting at an average accuracy of % when trained on and reaching % on the full dataset with an increase of %. The same observation can also be made for the IJB-B and IJB-C benchmarks (Table 2).
@@
The impact of the dataset depth (number of images) can be observed by comparing the achieved results of models trained/fine-tuned on K () and K (S) (). One can observe that increasing the dataset depth leads to higher verification accuracies in most settings.
@@ 5.3 Large-scale Fine-tuning
| Method | Backbone | Train data | LFW | CFP-FP | AgeDB30 | CALFW | CPLFW | Avg. | IJB-B | IJB-C | ||||
@@
| Method | Train data | LFW | CFP-FP | AgeDB30 | CALFW | CPLFW | Avg. | IJB-B | IJB-C |
| CosFace [60, 75] (CVPR 2018) | MS1MV2 | 99.81 | 98.18 | 98.34 | 96.18 | 92.76 | 97.05 | - | 96.01 |
| ArcFace [20, 75] (CVPR2019) | MS1MV2 | 99.78 | 98.54 | 98.21 | 96.05 | 92.93 | 97.10 | - | 96.03 |
| CurricularFace [32, 75] (CVPR2020) | MS1MV2 | 99.83 | 98.67 | 98.32 | 96.28 | 93.05 | 97.23 | - | 96.21 |
| MagFace [45] (CVPR2021) | MS1MV2 | 99.83 | 98.46 | 98.17 | 96.15 | 92.87 | 97.09 | 94.51 | 95.97 |
| ElasticFace-Arc [5] (CVPRW2022) | MS1MV2 | 99.80 | 98.67 | 98.35 | 96.17 | 93.27 | 97.25 | 95.22 | 96.49 |
| AdaFace [34] (CVPR2022) | MS1MV2 | 99.82 | 98.49 | 98.05 | 96.08 | 93.53 | 97.19 | 95.67 | 96.89 |
| CosFace [60, 75] (CVPR 2018) | WebFace4M | 99.80 | 99.25 | 97.45 | 95.95 | 94.40 | 97.37 | - | 96.86 |
| ArcFace [20, 75] (CVPR2019) | WebFace4M | 99.85 | 99.04 | 97.82 | 95.93 | 94.31 | 97.39 | - | 96.77 |
| CurricularFace [32, 75] (CVPR2020) | WebFace4M | 99.83 | 99.11 | 97.83 | 96.03 | 94.21 | 97.40 | - | 97.02 |
| AdaFace [34] (CVPR2022) | WebFace4M | 99.80 | 99.17 | 97.90 | 96.05 | 94.63 | 97.51 | 96.03 | 97.39 |
Table 3 presents the achieved recognition performances by models trained from scratch (baseline) and fine-tuned (CLIP and DINOV2) on a relatively small dataset, CASIA-WebFace (M images) and larger datasets, MS1MV2 (M images) and WebFace4M (M images). All results are reported using large (ViT-L) and small (ViT-S of DINOv2 and baseline and ViT-B of CLIP) architectures. We made the following observation from the reported results in Table 3:
Using CASIA-WebFace (M images), fine-tuning DINOv2 (ViT-S and ViT-L architectures) and CLIP (ViT-B and ViT-L architectures) led to higher verification accuracies, in comparison to the case where ViT-S and ViT-L are trained from scratch on the same data. For example, the average verification accuracies on the small benchmarks, LFW, CFP-FP, AgeDB30, CALFW, and CPLFW, was % by baseline (ViT-L) which is lower than % and % achieved by the fine-tuned DINOv2 and CLIP, respectively.
The models that are trained/fine-tuned on large datasets, MS1MV2 and WebFace4M, achieved higher verification accuracies than the one trained on the relatively smaller dataset, CASIA-WebFace.
Using small architectures (ViT-S and ViT-B) and large-scale training datasets (MS1MV2 or WebFace4M), the model trained from scratch achieved higher verification accuracies than the fine-tuned DINOv2 and CLIP. This observation can be noticed from the achieved results on small benchmarks as well as on large-scale, IJB-B and IJB-C, benchmarks.
Using large architectures (ViT-L) and large-scale training/fine-tuning datasets (MS1MV2 or WebFace4M), the fine-tuned DINOv2 and CLIP achieved slightly higher recognition performance than the models trained from scratch on most of the evaluation benchmarks.
To conclude, for the case where only a relatively small training dataset is accessible, fine-tuning pre-trained foundation models can achieve higher recognition accuracy than training the same model architecture from scratch. In the case that one has access to large training datasets, training a model from scratch for FR can achieve very competitive results to fine-tuning foundation models. However, this, next to the technical and legal limitations of collecting or maintaining the data, comes with a high training time cost. For example, training ViT-L from scratch on MS1MV2 using the settings described in Section 4 requires around GPU hours, in comparison to around GPU hours for fine-tuning DINOv2 with LoRA on Nvidia A100 SXM4 GB GPUs. Additionally, to put the performances achieved by FRoundation in the context of some of the major works in FR, we present in Table 4 results of the most influential works in the field [60, 20, 45, 5, 32, 34]. However, this comparison is not direct, as such works commonly use the ResNet-100 [29] architecture and the MS1MV2 or WebFace4M dataset [20, 75] for training. The various solutions outperform the trained from scratch models presented in Table 3. However, the results achieved by our FRoundation enhance the performance to a very close level to the works of the solutions presented in table 4.
| Method | Backbone | Train data | LFW | CFP-FP | AgeDB30 | CALFW | CPLFW | Avg. | IJB-B | IJB-C | ||||
| Baseline | ViT-S | SFace2 | 92.70 | 69.84 | 69.42 | 78.77 | 65.87 | 75.32 | 13.51 | 4.30 | 1.79 | 11.71 | 3.72 | 1.47 |
| IDiff-Face | 96.35 | 77.50 | 79.48 | 85.78 | 75.57 | 82.93 | 60.30 | 33.77 | 12.55 | 58.33 | 30.08 | 11.61 | ||
| ViT-L | SFace2 | 91.37 | 71.56 | 68.05 | 75.83 | 68.82 | 75.13 | 40.94 | 18.65 | 6.62 | 36.31 | 14.53 | 5.34 | |
| IDiff-Face | 96.13 | 76.56 | 77.43 | 85.25 | 75.83 | 82.24 | 56.93 | 26.14 | 6.07 | 54.30 | 21.20 | 4.20 | ||
| DINOv2 | ViT-S | - | 78.73 | 71.15 | 54.70 | 59.47 | 59.48 | 64.70 | 14.66 | 5.90 | 2.33 | 18.10 | 7.44 | 2.87 |
| SFace2 | 93.57 | 76.97 | 73.70 | 78.95 | 71.85 | 79.01 | 58.14 | 39.84 | 22.77 | 60.12 | 43.23 | 28.00 | ||
| IDiff-Face | 96.13 | 80.54 | 80.13 | 86.12 | 74.32 | 83.45 | 64.86 | 42.01 | 15.99 | 67.23 | 47.80 | 23.58 | ||
| ViT-L | - | 80.37 | 71.97 | 55.25 | 60.52 | 64.33 | 66.49 | 17.18 | 6.44 | 2.52 | 20.36 | 7.84 | 2.77 | |
| SFace2 | 95.55 | 84.58 | 77.73 | 82.48 | 78.10 | 83.69 | 68.33 | 48.92 | 28.30 | 71.37 | 53.32 | 34.94 | ||
| IDiff-Face | 97.77 | 86.94 | 86.67 | 90.28 | 81.73 | 88.68 | 75.11 | 55.75 | 32.80 | 78.75 | 62.17 | 40.94 | ||
| CLIP | ViT-B | - | 93.33 | 88.86 | 74.67 | 77.13 | 79.23 | 82.64 | 49.19 | 27.79 | 11.93 | 52.21 | 32.40 | 16.34 |
| SFace2 | 93.98 | 86.56 | 76.77 | 81.95 | 78.45 | 83.54 | 66.07 | 39.93 | 14.12 | 69.44 | 47.24 | 23.13 | ||
| IDiff-Face | 97.58 | 87.91 | 80.78 | 88.17 | 80.48 | 86.98 | 74.47 | 54.15 | 26.27 | 77.54 | 61.01 | 37.43 | ||
| ViT-L | - | 95.90 | 90.66 | 79.82 | 83.10 | 82.73 | 86.44 | 62.72 | 40.90 | 20.52 | 64.74 | 44.69 | 25.23 | |
| SFace2 | 97.88 | 89.87 | 79.03 | 86.50 | 83.88 | 87.43 | 75.31 | 54.86 | 25.14 | 78.64 | 60.81 | 37.73 | ||
| IDiff-Face | 98.75 | 91.40 | 86.50 | 90.78 | 84.87 | 90.46 | 82.92 | 64.56 | 34.82 | 86.07 | 71.77 | 48.18 | ||
5.4 Learning from Synthetic Data
Table 5 presents the verification accuracies obtained by models trained from scratch (baseline) on IDiff-Face [6] and SFace2 [7] alongside the considered foundation models, DINOv2 and CLIP, which were fine-tuned using LoRA on IDiff-Face and SFace2. All results are reported using the experimental setups presented in Section 4. From the results presented in Table 5, we draw the following insights:
The recognition performance of CLIP and DINOv2 improves when fine-tuning with any of the considered synthetic datasets compared to using the pre-trained models alone, highlighting the significant potential of synthetic datasets in enhancing model performance.
The models trained/fine-tuned on the IDiff-Face dataset generally deliver higher recognition performance across various benchmarks compared to those trained on the SFace2 dataset. For example, the ViT-L model of CLIP achieves higher verification accuracy on all evaluation benchmarks when fine-tuned on IDiff-Face, compared to when it is fine-tuned on SFace2. On the small benchmarks, the model achieved an average accuracy of 90.46% when fine-tuned on IDiff-Face, compared to 87.43% when fine-tuned on SFace2. A similar trend is observed on the large benchmarks, where the model fine-tuned on IDiff-Face achieved accuracies of 64.56% and 71.77% on IJB-B and IJB-C at a FAR of 1e-4, respectively, while the model fine-tuned on SFace2 achieved accuracies of 54.86% and 60.81% on the same benchmarks.
Fine-tuning the larger ViT models of pre-trained DINOv2 and CLIP consistently led to higher verification accuracies on most benchmarks compared to their smaller counterparts. This is not the case for the baseline models, where both the ViT-S and ViT-L demonstrate competitive performance across various benchmarks. This can be attributed to the fact that larger models are more prone to overfitting when trained on smaller-scale datasets [23].
This investigation highlights the potential of synthetic data in enhancing the performance of foundation models for FR, demonstrating superior results compared to both pre-trained and baseline models. Furthermore, it encourages further exploration of additional synthetic datasets for FR, such as DCFace [35] and the more recent ID3 [65].